OpenStudy (anonymous):

what is 12/28 in simplest form

OpenStudy (michele_laino):

@sijieh111 please calculate the maximum common divisor between 12 and 28

OpenStudy (anonymous):

\[\frac{ 12 }{ 28 }\] both are divisible by 4 so \[\frac{ \frac{ 12}{ 4 } }{ \frac{ 28 }{ 4 } }\]

OpenStudy (michele_laino):

do you know how to calculate the maximum common divisor between two numbers?